# Towards Open Federated Learning Platforms: Survey and Vision from Technical and Legal Perspectives

### 18 Jan 2024 v2

@@ Abstract.
Traditional Federated Learning (FL) follows a server-domincated cooperation paradigm which narrows the application scenarios of FL and decreases the enthusiasm of data holders to participate. To fully unleash the potential of FL, we advocate rethinking the design of current FL frameworks and extending it to a more generalized concept: Open Federated Learning Platforms. We propose two reciprocal cooperation frameworks for FL to achieve this: query-based FL and contract-based FL. In this survey, we conduct a comprehensive review of the feasibility of constructing an open FL platform from both technical and legal perspectives. We begin by reviewing the definition of FL and summarizing its inherent limitations, including server-client coupling, low model reusability, and non-public. In the query-based FL platform, which is an open model sharing and reusing platform empowered by the community for model mining, we explore a wide range of valuable topics, including the availability of up-to-date model repositories for model querying, legal compliance analysis between different model licenses, and copyright issues and intellectual property protection in model reusing. In particular, we introduce a novel taxonomy to streamline the analysis of model license compatibility in FL studies that involve batch model reusing methods, including combination, amalgamation, distillation, and generation. This taxonomy provides a systematic framework for identifying the corresponding clauses of licenses and facilitates the identification of potential legal implications and restrictions when reusing models. Through this survey, we uncover the the current dilemmas faced by FL and advocate for the development of sustainable open FL platforms. We aim to provide guidance for establishing such platforms in the future, while identifying potential problems and challenges that need to be addressed.
1. Introduction
@@
In recent years, the barriers to the development of Artificial Intelligence (AI) have been broken down with the rapid progress of ABC technologies in computing: AI, Big Data, and Cloud Computing, as well as the emergence of cost-effective specialized hardware (Sze et al., 2017) and software (Jia et al., 2014). This has led to the world entering the third wave of AI development: Deep Learning (LeCun et al., 2015). The success of current data-driven AI relies on massive amounts of training data and follows a gather-and-analyze paradigm (Whang et al., 2023), which confronts with challenges of complying with rigorous data protection regulations such as OECD Privacy Guidelines (Tene, 2011) and General and Data Protection Regulation (GDPR) (Voigt and Von dem Bussche, 2017). So although data-centric AI is currently mainstream paradigm, Federated Learning (Li et al., 2020e), a novel model-centric distributed collaborative training framework, is gaining popularity in both academia and industry for its advantages in complying with privacy regulations (Truong et al., 2021).
@@
According to the definitions of IEEE Standard for Federated Machine Learning (FML, aka FL) (Society, 2021), FL is a framework or system that enables multiple participants to collaboratively build and use machine learning models without disclosing the raw and private data owned by the participants while achieving good performance. For example, a typical workflow of FL systems is that the entity with modeling demand (aka FL server) first deploys the FL services and initializes the model training task, and then distributing this task to participants with training data (aka FL clients) for modeling (Bonawitz et al., 2019). Based on this workflow pattern, many FL frameworks have been derived with specialized improvements in communication (Konečnỳ et al., 2016; McMahan et al., 2017; Xu et al., 2021b), optimizaiton (Li et al., 2020f; Karimireddy et al., 2020; Li et al., 2021b), robustness (Duan et al., 2020; Sattler et al., 2019; Li et al., 2022d) and privacy (Bonawitz et al., 2017; Geyer et al., 2017; Cheng et al., 2021). While these fascinating improvements greatly enhance the utility of FL, they all follow a task-based interaction paradigm, in which an FL server dominates the cooperation between FL participants. In this narrow interpretation of FL, the data owner is treated more like a worker than a collaborator and performs training primarily for the benefit of the server’s goals. Due to the above defects, clients have little enthusiasm to participate, and the potential for redundant training also leads to low model reusability, further diminishing the efficiency of the FL systems. This explains why current FL frameworks are more akin to private distributed modeling services rather than sustainable and privacy-preserving modeling platforms for everyone as expected.
@@
Figure 1. A schematic diagram of three cooperation frameworks of FL. (a) (b) are the proposed open FL platforms, (c) is the traditional FL platform. Four colors correspond to four roles in (Society, 2021), and colors with grid lines indicate non-essential roles.
[…]
@@
In this paper, we try to answer the question: Can we establish a sustainable open FL platform based on a novel reciprocal cooperation framework? Obviously, to answer this quesion, it is insufficient simply study the basic concepts of FL and investigate existing FL techniques. We also need to conduct a wide survey of potential techniques that can facilitate the construction of open FL platforms. To aid understanding, Fig. 1 provides a first glimpse of two novel FL cooperation frameworks we advocated:
@@
  • Query-based FL. It follows a loosely-coupled cooperation framework between entities (we use ”entities” instead of ”participants” to emphasizes equality), where any entity can freely upload their local models or retrieve models from an open repository named Model Community. There are many valuable challenges that can be explored, such as how to query for models, how to reuse the retrieved models or how to transfer knowledge from these models, how to ensure the legal compliance between different model licenses, how to protect the intelligent properties of released models (ref. Section 4).
@@
  • Contract-based FL. It follows a mutual choice cooperation framework, where each entity can deploy model training contracts with specialized requirements such as task modality, execution environment, model architecture and license. Meanwhile, entities holding data can choose whether to accept the contract. Research topics in this setting include ML subtasks design and monetization (ref. Section 5).
@@
It’s worth noting that the definitions of the four roles illustrated in Fig. 1 (i.e. model user, coordinator, data owner, auditor, ref. Section 3.1) are adopted for compatibility with the IEEE standard (Society, 2021), and our proposals are also within the standard definitions of FML. The diagram in Fig. 1(c) illustrates the workflow of traditional FL, where all FL clients are required to accept the training schedule from the FL server and perform multiple rounds of local training and model averaging until the global model converges. In contrast, the entities in query-based FL and contract-based FL are proactive in their participate. We believe that these reciprocal cooperation frameworks have the potential to expand the prevalence of open FL and establish FL ecosystems.
1.1. Our Contribution
@@
In contrast to previous surveys that primarily focused on the server-dominated cooperation framework in FL, our new survey explores the feasibility of reciprocal cooperation frameworks in FL. To the best of our knowledge, our work represents the first systematic survey in this area. The major contributions of this survey are as follows:
@@
  • We introduce the concept of open FL platforms by presenting two reciprocal cooperation frameworks, namely query-based FL and contract-based FL, along with an overview of their key features and properties.
@@
  • We propose a new taxonomy to streamline the legal compliance analysis in FL studies, which is also useful for quickly identifying suitable model reusing methods for open FL platforms. A comprehensive comparison of current FL studies based on this taxonomy is surveyed.
@@
  • We analyze the requirements for model protection in the context of query-based FL and identify applicable solutions from deep intellectual property protection technologies.
@@
The rest of this paper is organized as follows. We compare this survey to other related surveys and show our distinction in Section 2. In Section 3, we presente the overview and point the limitations of traditional FL. We comprehensively explore the feasibility and challenges of query-based FL in Section 4, which includes model query (Section 4.2), model license comparison (Section 4.3.1) and selection (Section 4.3.2), copyright issues (Section 4.3.3), and analysis of license conflicts in model reusing (Section 4.3.6). In Section 4.3.4, we present our taxonomy from a model reusing perspective, and in Section 4.3.5, we summarize FL studies based on this new taxonomy. The discussion on model intellectual property protection is presented in Section 4.4.
2. Related Surveys
@@
Federated learning has become a buzzword in various fields, leading to the emergence of numerous FL studies. These works can be classified into three primary categories: FL systems design, FL applications and FL toolkits. Extensive surveys are available to summarized the advancement of federated learning, as shown in Table 1. The initial architectures and concepts for FL systems were summaried by Yang et al. (Yang et al., 2019). They categorized FL into horizontal FL, vertical FL and federated transfer learning based on the distribution characteristics of data, which are written in IEEE Standard 3652.1-2020 (Yang et al., 2021; Society, 2021). Following this, an increasing number of surveys have emerged focusing on enhancing FL system design (Li et al., 2020e; Aledhari et al., 2020; Kairouz et al., 2021; Zhang et al., 2021b; Li et al., 2021c). From the algorithmic perspective, personlized FL (Kulkarni et al., 2020; Tan et al., 2022b) aims to learn personlized models for each client to address the challenge of statistical heterogeneity (Ma et al., 2022). Besides, the privacy-perserving computing platforms and model aggregation protocols for FL systems also been widely studied and sumaried by (Liu et al., 2022; El Ouadrhiri and Abdelhadi, 2022; Yin et al., 2021; Lyu et al., 2020). Furthermore, many advanced FL architectures had been proposed, such as asynchronous (Xu et al., 2021b), decentralized and blockchain-based FL frameworks (Nguyen et al., 2021; Qu et al., 2022b; Zhu et al., 2022). Given that federated learning technologies enable collaboration among distributed participants in model training and decision-making, this capability holds great promise in a wide range of application scenarios. For instance, multiple geogrphically distributed medical insitutions can enhace medication recommendation, drug-drug interaction prediction and medical image analysis in a collaborative manner without exchanging any sensitive data (Xu et al., 2021a; Pfitzner et al., 2021; Antunes et al., 2022; Rieke et al., 2020). The massive real-time data generated by IoT devices in smart cities (Zhang et al., 2022b; Ramu et al., 2022), industries (Boopalan et al., 2022), vehicles (Du et al., 2020) has also sparked interest in exploring how FL technology can be used to deliver more advanced services such as intrusion detection, anomaly detection, fraud detection and network load prediction (Agrawal et al., 2022; Alazab et al., 2021; Ghimire and Rawat, 2022).
@@
As summarized in Table 1, most surveys extensively discuss the challenges of efficiency, heterogeneity, privacy in FL systems design, with the surveys from blockchain fileds offering the most comprehensive review. However, except for a few blockchain-based FL studies, most of the listed surveys just present the same story from slightly different angles and backgrounds, i.e. a server sets the model training task and delegate it to data holders to complete. This server-dominated cooperation framework is a narrow implementation of the FL systems. Therefore, this survey aim to fill the gap by investigating and surveying the associated tenchnologies that support more open and inclusive cooperation frameworks in FL systems, where all entities, whether they own the data or not, can benefit from it. The challenges investigated in this survey are not listed in the Table 1, to the best of our knowledge, this is the first survey that focuses on the cooperation frameworks of FL.
@@
Distinction of Our Survey. This survey focuses on exploring the innovative cooperation frameworks in FL, which will involve some FL concepts such as decentralized FL, blockchain-based FL, few-shot FL, ML related platforms and services but goes beyond them. In the following sections, we will distinguish our survey by highlingting the similarities and differences between these related concepts.
Table 1. Summary of existing FL surveys, SYS denotes FL Systems Design, APP denotes FL Applications, SDC denotes Server-Dominated Cooperation frameworks.
Challenges Contents
Scenarios/Tasks FL Surveys Efficiency Heterogeneity Privacy Incentive Decentralized SYS APP SDC
General Yang et al. (Yang et al., 2019)
Li et al. 2020 (Li et al., 2020e)
Zhang 2021et al. (Zhang et al., 2021b)
Gupta et al. (Gupta and Alam, 2022)
Xu et al. (Xu et al., 2021b)
Li et al. 2021 (Li et al., 2021c)
El et al. (El Ouadrhiri and Abdelhadi, 2022)
Kulkarni et al. (Kulkarni et al., 2020)
Liu et al.(Liu et al., 2022)
Tan et al. (Tan et al., 2022b)
Zhu et al. 2021 (Zhu et al., 2021b)
Ma et al. (Ma et al., 2022)
Aledhari et al. (Aledhari et al., 2020)
Kairouz et al. (Kairouz et al., 2021)
AbdulRahman et al. (AbdulRahman et al., 2020b)
Lim et al. (Lim et al., 2020)
Healthcare Xu et al. (Xu et al., 2021a)
Pfitzner et al.(Pfitzner et al., 2021)
Antunes et al. (Antunes et al., 2022)
Rieke et al. (Rieke et al., 2020)
IoT Zhang 2022et al. (Zhang et al., 2022b)
Boopalan et al. (Boopalan et al., 2022)
Ramu et al. (Ramu et al., 2022)
Du et al. (Du et al., 2020)
Cybersecurity Agrawal et al. (Agrawal et al., 2022)
Alazab et al. (Alazab et al., 2021)
Ghimire et al. (Ghimire and Rawat, 2022)
Blockchain Nguyen et al. (Nguyen et al., 2021)
Qu et al. (Qu et al., 2022b)
Zhu et al. 2022 (Zhu et al., 2022)
2.1. FL Systems
Federated learning, with its nature advantages in privacy-preserving decision sharing, has garnered significant attention in both industry and academia, leading to the rapid development of federated learning systems. The earliest attempt at the large-scale FL system was by Google, where FL was used to improve next-word prediction (Hard et al., 2018) and query suggestion (Yang et al., 2018) for Gboard applications. Subsequently, many novel FL systems have emerged to adapt to diverse federated training scenarios, such as Horizontal FL (e.g., TFF (Abadi et al., 2016), FedLab (Zeng et al., 2021), Felicitas (Zhang et al., 2022f), IBM FL (Ludwig et al., 2020), OpenFL (Foley et al., 2022)), Vertical FL (Wu et al., 2022a) or both (e.g., FATE (Liu et al., 2021b), FedML (He et al., 2020b), PaddleFL (Ma et al., 2019), Flower (Beutel et al., 2020), FedTree (Li et al., 2022b), NVFLARE (Roth et al., 2022)). Despite these frameworks covering a wide range of application scenarios, they all follow the server-dominated cooperation mechanism. This business model restricts FL to function as a collaborative modeling software, rather than an open platform that provides federated training services to the public.
Unlike the FL systems mentioned above, PySyft (Ziller et al., 2021) developed by OpenMined depicts a novel FL cooperation frameworks which is closely realted to our focus. PySyft encourages data owners to share their data on a private domain server, which provides data management and privacy controls, as well as limited machine learning analysis APIs for third-party data scientists. Besides, a public network server will provide connections between data owners and data scientist, enabling datasets search and discovery for platform users. Recently, a new FL platform named PySyTFF [^1] https://blog.openmined.org/announcing-proof-of-concept-support-for-tff-in-pysyft-0-7/ was announced. It integrates TFF and PySyft, allowing data scientists to train models under the coordination of TFF and the datasets provided by PySyft domain servers. However, even with inference controls of datasets, there is still a high security risk associated with exposing access to sensitive data on the Internet (Gamundani and Nekare, 2018). To preserve the privacy advantages of FL, in this survey, we investigate open and data-free FL platforms under the scope of model-centric ML (Lou et al., 2020). In such FL platforms, every user is free to collaborate on the training of machine learning models while privacy is protected.
2.2. As-a-Service Business Model
In the current context of Software-as-a-Service (SaaS) (Brereton et al., 1999), there are several as-a-service cloud computing frameworks that encapsulate ML tasks as services and provides unified APIs for upper layer applications. For example, Model-as-a-Service (MaaS) (Geller and Turner, 2007; Roman et al., 2009; Zou et al., 2012; Liu et al., 2021c; Sun et al., 2022b) and Machine-Learning-as-a-Service (MLaaS) (Ribeiro et al., 2015; Hanzlik et al., 2021; Hesamifard et al., 2018; Li et al., 2017a; Kourtellis et al., 2020) encapsulate model execution and model development as services. The original concept of MaaS (Geller and Turner, 2007; Roman et al., 2009) was to provide re-usable and fine-grained user interfaces and visualization tools of domain-specific models (e.g wealther model, oil spill detection model) for environmental decision support systems. Subsequently, this concept has been extended to the field of recommendation systems (Zou et al., 2012) and deep learning based systems (Liu et al., 2021c; Sun et al., 2022b). However, in contrast to the focus of this survey, the aforementioned MaaS framework does not involve any user collaboration but solely provides model inference APIs to users.
As the architectures of deep neural networks (DNNs) become increasingly complex, training and maintaining DNNs become more and more challenging (Han et al., 2021). To address this issue, cloud service providers have introduced MLaaS, which offers an integrated development environment as a service for constructing and operationalizing ML workflows, aiming to reduce the computational resources required. MLaaS enables users to upload their data for training (Ribeiro et al., 2015; Zhao et al., 2021; Hesamifard et al., 2018) or inference (Hanzlik et al., 2021), freeing them from the responsibility of managing hardware resources and implementation. Most MLaaS providers adopt a pay-by-query business model, such as Google Vertex AI [^2] https://cloud.google.com/vertex-ai, Microsoft Azure Machine Learning [^3] https://azure.microsoft.com/products/machine-learning/ and ChatGPT [^4] https://chat.openai.com/chat. However, privacy protection can be compromised when users upload data to perform inference and training in the cloud. Moverover, under this model, users are not given the ability to contribute their own models to the repository or collaborate with others to enhance the diversity of available models. While there are some ongoing efforts to offer privacy-preserving MLaaS services using techniques such as Trusted Execution Environment (TEE) (Hanzlik et al., 2021; McKeen et al., 2016) and Homomorphic Encryption (Hesamifard et al., 2018; Gentry, 2009), it is worth noting that our focus is not solely on privacy.
Recently, Kourtellis et al. (Kourtellis et al., 2020) propose Federated Learning as a Service (FLaaS) that provides high-level and extensible APIs aim to enabling third-party applications to build collaborative, decentralized, privacy-preserving ML models. Jiang te al. (Jiang et al., 2022) propose an open FL ecosystem for mobile devices, which shares a similar concept to FLaaS. However, those approach also follow the traditional server-dominated cooperation framework, which falls under the scope of previous FL surveys(Yang et al., 2019; Li et al., 2020e; Kairouz et al., 2021).
2.3. Dcentralized FL
Decentralized FL (Lalitha et al., 2018; Kalra et al., 2023; Marfoq et al., 2020; Hu et al., 2019; Sun et al., 2022a), a novel server-less paradigm of FL, emphasizes the advantages of employing a peer-to-peer model delivery and aggregation network that is free from the dependencies of a central trusted server. Instead of solely communicating with a central server, participants can fully leverage the network bandwidth by utilizing the network connections between them. For example, Lalitha et al. (Lalitha et al., 2018) proposed exchange and merge of posterior distribution among neighboring users to collaboratively estimate the global optimal parameter. Similar to the local training of FedAvg, DFedAvgM (Sun et al., 2022a) also suggests that each client communicates with its neighbors after multiple training iterations to improve the convergence rate of training. On the other hand, ProxyFL (Kalra et al., 2023) improves the privacy of neighbor-wide model sharing by sharing a proxy model through knowledge distillation (Hinton et al., 2014). However, the major bottleneck lies in the high communication cost of sharing the model with all neighbors in a fully connected network.
To address this issue, Marfoq et al. (Marfoq et al., 2020) proposed improving the efficiency of model sharing by selecting a connected subgraph. Another approach is the use of a gossip-based approach, where model parameters or segments of model parameters are randomly shared with peer neighbors (Hegedűs et al., 2021; Hu et al., 2019). Despite the advantages brought by the decentralized design, the training procedure of these frameworks also follows a preset learning task and lacks sustainable cooperation, resulting in a non-public and low reusability FL platform similar to centralized FL. In fact, our vision for open FL platforms is to extend the FAIR principles (Wilkinson et al., 2016) for scientific data to the context of machine learning. We believe that all dedicated models in these platforms should adhere to the principles of being Findable, Accessible, Interoperable, and Reusable.
3. Basic Concepts of Federated Learning
3.1. Definition

### 29 Feb 2024 v3

@@ Abstract
Traditional Federated Learning (FL) follows a server-dominated cooperation paradigm which narrows the application scenarios of FL and decreases the enthusiasm of data holders to participate. To fully unleash the potential of FL, we advocate rethinking the design of current FL frameworks and extending it to a more generalized concept: Open Federated Learning Platforms, positioned as a crowdsourcing collaborative machine learning infrastructure for all Internet users. We propose two reciprocal cooperation frameworks to achieve this: query-based FL and contract-based FL. In this survey, we conduct a comprehensive review of the feasibility of constructing open FL platforms from both technical and legal perspectives. We begin by reviewing the definition of FL and summarizing its inherent limitations, including server-client coupling, low model reusability, and non-public. In particular, we introduce a novel taxonomy to streamline the analysis of model license compatibility in FL studies that involve batch model reusing methods, including combination, amalgamation, distillation, and generation. This taxonomy provides a feasible solution for identifying the corresponding license clauses and facilitates the analysis of potential legal implications and restrictions when reusing models. Through this survey, we uncover the current dilemmas faced by FL and advocate for the development of sustainable open FL platforms. We aim to provide guidance for establishing such platforms in the future while identifying potential limitations that need to be addressed.
Index Terms:
Federated Learning, AI Licensing, Collaborative Machine Learning, Model Mining
I Introduction
@@
The success of current data-driven AI relies on massive amounts of training data and follows a gather-and-analyze paradigm [1], which confronts with challenges of complying with rigorous data protection regulations such as OECD Privacy Guidelines [2] and GDPR [3]. Although data-centric AI is currently mainstream paradigm, Federated Learning [4], a novel distributed collaborative training framework, is gaining popularity in both academia and industry for its advantages in complying with privacy regulations [5].
@@
A typical workflow of FL systems is presented in Fig.1(a), where the entity with a modeling demand (FL server) first deploys the FL services, initializes the training task, and then distributes this task to participants with training data (FL clients) for modeling[6]. Based on this workflow pattern, FL frameworks have been derived with specialized improvements in communication [7, 8, 9], optimizaiton [10, 11, 12], robustness [13, 14, 15] and privacy [16, 17, 18]. While these fascinating improvements greatly enhance the utility of FL, they all follow a task-based interaction paradigm, in which an FL server dominates the cooperation. In this narrow interpretation of FL, the data owner is treated more like a worker than a collaborator and performs training primarily for the benefit of the server’s goals. Due to the above defects, FL clients have little enthusiasm to participate, and the potential for redundant training also leads to low model reusability, further diminishing the efficiency of the FL systems. This explains why current FL frameworks are more akin to private distributed modeling services rather than open and sustainable modeling platforms that every user can access and benefit from, addressing many data silo applications.
@@
Figure 1: A schematic diagram of three cooperation frameworks of FL. (a) is the traditional FL platform, (b) (c) are the proposed open FL platforms. Four colors correspond to four roles in [19], and colors with grid lines indicate non-essential roles.
@@
In this paper, we try to answer the question: Can we establish a sustainable open FL platform based on a novel reciprocal cooperation framework? Obviously, to answer this quesion, it is insufficient to simply study the basic concepts of FL and investigate existing FL techniques. We also need to conduct a wide survey of potential techniques that can facilitate the construction of open FL platforms. To aid understanding, Fig. 1(b)(c) provide a first glimpse of two novel FL cooperation frameworks that we advocated:
@@
  • Query-based FL. It follows a loosely-coupled cooperation framework between entities (we use "entities" instead of "clients" to emphasizes equality), where any entity can freely upload their local models or query models from an open repository named Model Community. There are many valuable challenges that can be explored, such as how to query for models, how to reuse the retrieved models or how to transfer knowledge from these models, how to ensure the legal compliance between different model licenses, and how to protect the intelligent properties of released models (ref. §IV, §V and ‡II, ‡III, ‡IV, ‡V, where § denotes the main section and ‡ denotes the appendix section).
@@
  • Contract-based FL. It follows a mutual choice cooperation framework, where each entity can deploy model training contracts with specialized requirements such as task modality, execution environment, model architecture and license. Meanwhile, entities holding data can choose whether to accept the contract. We show the research topics in this setting include ML subtasks design and monetization (ref. §VII and ‡VI).
@@
It is worth noting that the definitions of the four roles illustrated in Fig. 1 (i.e., model user, coordinator, data owner, auditor, ref. §III-A) are adopted for compatibility with the IEEE standard [19], and our proposals are also within the standard definitions of FL. The diagram in Fig. 1(a) illustrates the workflow of traditional FL, where all FL clients are required to accept the training schedule from the FL server and perform multiple rounds of local training and model averaging until the global model converges. In contrast, the entities in query-based FL and contract-based FL are proactive in their participation. We believe that these reciprocal cooperation frameworks have the potential to expand the prevalence of open FL and establish FL ecosystems.
I-A Our Contribution
@@
In contrast to previous surveys that primarily focused on the server-dominated cooperation framework in FL (ref. §II), our new survey explores the feasibility of reciprocal cooperation frameworks in FL. To the best of our knowledge, our work represents the first systematic survey in this area. The major contributions of this survey are as follows:
@@
  • We are the first to introduce the concept of open FL platforms by presenting two cooperation frameworks, namely query-based FL and contract-based FL, along with an overview of their key features and properties.
@@
  • We propose a taxonomy to streamline the legal compliance analysis in ML, which is also useful for quickly identifying suitable model reusing conceptss for open FL platforms. A comprehensive comparison of current FL studies based on this taxonomy is surveyed.
@@
  • We analyze the requirements for model protection in the context of query-based FL and identify applicable solutions from deep Intellectual Property (IP) protection. We also introduce the concept of designing ML microtasks for query-based FL.
@@
The structure of this paper is shown in Fig. 2. We compare this survey to other related surveys and show our distinction in §II. In §III, we present the overview and point the limitations of traditional FL. We comprehensively explore the feasibility and challenges of query-based FL in §IV, which includes model query (§IV-B), model license comparison (§IV-C1) and license selection (§IV-C2). In §V, we present our taxonomy from a model reusing perspective. We introduce how to design ML microtasks in contract-based FL in §VII and we conclude this paper in §VIII. Due to page limits, we defer some related discussions to the appendix and we use ‡ to remind you that detailed analysis can be found in the appendix.
§I. Introduction §II. Related Surveys §III. Basic FL Concepts §IV. Query-based FL §V. Batch Model Reuse Mechanisms §VI. Remaining Topics §VII. Contract-based FL §VIII. Conclusion §I-A. Our Contribution ‡I. Related Topics As-a-Service Dcentralized FL FL Systems §III-A. Definition §III-B. Limitations of FL §IV-A. Overview §IV-B. How to Query §IV-C. Legal Issues ‡II. Filter Conditions ‡III. License Choosing Copy-rightability Datasets, Software, Models Licenses Combination, Amalgam…, Distillation, Generation §IV-B. Hybrid Model Reusing… ‡IV-B. Model Reusing vs. Model Licenses ‡IV-A. Hybrid Model Reusing in FL §VI-A. Model Protection §VI-B. Limitations of Query-based FL ‡V. How to Protect Models §VII-A. Overiew ‡VI. Decentralization and Monetization in Contract-based FL §VII-B. How to Design ML Microtasks §VII-C. Limitations of Contract-based FL
Figure 2: Article Structure. §: Main SectionSubsection, ‡: Appendix.
II Related Surveys
@@
Federated learning has become a buzzword in various fields, leading to the emergence of numerous FL studies. These works can be classified into three primary categories: FL systems design, FL applications and FL toolkits. The initial architectures and concepts for FL systems were summaried by Yang et al. [20]. They categorized FL into horizontal FL, vertical FL and federated transfer learning based on the distribution characteristics of data, which are written in IEEE Standard 3652.1-2020 [21, 19]. Following this, several surveys have emerged with a focus on enhancing FL system [4, 22, 23, 24, 25]. From the algorithmic perspective, personlized FL [26, 27] aims to learn personlized models for each client to address the challenge of statistical heterogeneity [28]. Meanwhile, the privacy-perserving computing platforms and model aggregation protocols for FL also been widely studied and summarized by [29, 30, 31, 32]. Furthermore, many advanced FL architectures had been proposed, such as asynchronous [9], decentralized and blockchain-based FL frameworks [33, 34, 35].
@@
Currently, most surveys extensively discuss the challenges of efficiency, heterogeneity, privacy in FL systems design, while the surveys from blockchain fields offer the most comprehensive review‡I. However, except for a few blockchain-based FL studies, most of the listed surveys just present the same story from different angles and backgrounds, i.e., a server sets the model training task and delegates it to data holders to complete. This server-dominated cooperation framework is a narrow implementation of FL systems. Therefore, this survey aims to fill the gap by investigating and surveying the associated tenchnologies that support more open and inclusive cooperation frameworks in FL systems, where all entities, whether they own the data or not, can benefit from it.
@@
Distinction of Our Survey. This survey focuses on exploring the innovative FL cooperation frameworks, which involves some FL concepts such as decentralized FL, blockchain-based FL, few-shot FL, ML related platforms and services but goes beyond them. To the best of our knowledge, this is the first survey that focuses on the cooperation frameworks of FL‡I.
III Basic Concepts of Federated Learning
III-A Definition